OpenStudy (anonymous):

According to the general equation for conditional probability, if P(A^B)= 2/3 and P(B)= 3/4, what is P(A|B)? A. 15/16 B. 24/25 C. 8/9 D. 35/36

jimthompson5910 (jim_thompson5910):

hint: \[\Large P(A|B) = \frac{P(A \cap B)}{P(B)}\]

OpenStudy (anonymous):

Maybe I'm just an idiot? I'm still not getting any of the answers

jimthompson5910 (jim_thompson5910):

Do you have this so far? \[\Large P(A|B) = \frac{P(A \cap B)}{P(B)}\] \[\Large P(A|B) = \frac{2/3}{3/4}\]

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

when you divide 2/3 over 3/4, you're really multiplying 2/3 by 4/3 the second fraction flips and you multiply

OpenStudy (anonymous):

the correct answer was 8/9

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

since 2/3 times 4/3 = 8/9